Find the exact values of sin 2u, cos 2u, and tan 2u using the double-angle formulas.cos u = − 15/17, π/2 < u < π
Answer by Alan3354(69443)

Find the exact values of sin 2u, cos 2u, and tan 2u using the double-angle formulas.cos u = − 15/17, π/2 < u < π
--------------
Find sin(u):
sin = sqrt(1 - cos^2) = +8/17 (+ in Q2)
------
sin(2u) = 2*sin(u)*cos(u)
---
Then cos(2u) = sqrt(1 - sin^2(2u))
Then tan(2u) = sin(2u)/cos(2u)